# The Contiguous Art Gallery Problem is in $\Theta(n\log n)$

### v1

@@ 3.2  Lower bound
In Section˜10, we reduce from Set Disjointness, where we are given two sets $A,B\subset[0,n^{3}]$, each of size $n$, and must decide whether $A\cap B=\emptyset$. While in the wordRAM-model there is a deterministic $O(n\log\log n)$-time algorithm for solving this problem (via sorting [23]), in comparison-based models of computation such as the realRAM-model, it has a worst-case lower bound of $\Omega(n\log n)$, even if $A$ is already sorted [45]. Given $A,B\subset[1,n^{3}]$ with $A$ sorted, we can in linear time construct a simple polygon $P$ such that $A\cap B=\emptyset$ if and only if $P$ can be guarded with $2|B|+2$ guards (see Figure˜2).
@@ 4  A combinatorial classification of guards
We will define three collections of dominators: the good, the bad, and the ugly. We prove that all good guards are dominated by a good dominator, which will induce a set $X$ of $O(n)$ points along $\partial P$ such that there exists an optimal solution that includes a guard $(g,[x,v])$ for $x\in X$. We prove that all bad guards are dominated by either a bad or an ugly dominator. The flowchart in Figure 3 illustrates our approach. To this end, we first make some observations:
@@ 4.3  The next function and our high-level algorithm
Definition 8 (The next function [9, 38]).
@@ 5  Intermezzo: data structures and sliding sequences
Data Structure 1 ([22]).
@@
The data structure of [22] stores a family of shortest paths in $P$ as balanced binary trees. For any query points $s,t\in P$, the shortest path $S(s,t)$ can be expressed as $O(\log n)$ subtrees, together with $O(\log n)$ newly computed edges called bridges.
The second property follows from a standard adaptation of this data structure (see also [20]): for each pair of consecutive edges on a precomputed shortest path, determine whether the turn is left or right. For each node in the tree, store a symbol indicating whether all of its descendant edges are left-turning, right-turning, or mixed. For a query, we can determine for each of the $O(\log n)$ newly computed bridges whether the turn at its endpoints is left or right. Combined with the pre-stored symbols, this information suffices to answer the second query. ∎
Data Structure 2 ([24]).
@@
Throughout the paper, we assume that we have access to ˜1 and ˜2 for the given polygon. We repeatedly use the concept of a sliding sequence. A sliding sequence can be thought of as a pair of indices $(i,j)$ that can be updated by incrementing either $i$ or $j$. Formally we define in Definition˜1 a conforming sliding sequence used by almost all of our algorithms. For each $(i,j)$, we maintain the visibility core $\mathcal{E}[i-1,j+1]$ as a convex-hull data structure supporting emptiness, containment, extreme-point, and ray-shooting queries, and an additional angle query (defined later). For any linear-size sliding sequence $\{(i_{t},j_{t})\}$, Chan, Hershberger, and Pratt [13] give a data structure supporting such updates and queries in $O(\log n)$ time:
Lemma 9 (Paraphrased Lemma 1 in [13]).
@@
The above two lemmas imply the following, see also the remarks made in [13].
@@
Let $(i,j)$ be the dynamic index pair from the Sliding-Window data structure (3). We maintain $\mathcal{E}[i-1,j+1]$ and $\mathcal{E}[i-1,j+2]$ separately, subject to increments of $i$ and $j$ via Corollary˜11. This maintains the vertices of these convex areas as balanced binary trees where the left-to-right traversal of the leaves corresponds to a clockwise traversal. For a a convex polygon with $n$ vertices, whose vertices are stored in a balanced tree according to their cyclic order, the containment, ray shooting, and extreme-point queries can be answered in $O(\log n)$ time [18]. We can thus immediately support all queries except for angle queries.
@@ 6.3  A linear-size set $\mathcal{B}_{\sigma}$ that contains all reduced good dominators in $\mathcal{B}$
We store the active points in the data structure of Brodal and Jacob [11], which supports insertions and deletions in $O(\log n)$ time and and extreme-point queries in $O(\log n)$ time. Each time we add a vertex to the active set, we perform an insertion. When $j^{\prime}$ is incremented, we let $H$ be the half-plane left of $j^{\prime}+1,j^{\prime}+2]$ and iteratively perform extreme-point queries in the direction of the outward normal of $H$. If a query returns a vertex $x$ outside $H$, then we set $r_{x}=[j^{\prime}+1,j^{\prime}+2]$ and remove $x$ from the active set, charging $O(\log n)$ time for both the query and the update. The first time a query returns a vertex inside $H$, all remaining active points lie in $H$, and the process terminates.
@@
To construct the corresponding querying data structure, we loop over all $(i,j)\in\sigma$ in order, which discretizes time $t\in[0,|\sigma|]$ where at time $t$ our loop is at $(i_{t},j_{t})\in\sigma$. For each $g\in V_{\sigma}$ there is a unique time interval $t_{1},t_{2}]$ where for all $t\in[t_{1},t_{2}]$, $g$ is a vertex of $\mathcal{E}[i_{t}-1,j_{t}+1]$. We loop over all $t$ and use the first-in-first-out convex hull data structure from Corollary˜11 to maintain $\mathcal{E}[i_{t}-1,j_{t}+1]$ explicitly in $O(n\log n)$ time and $O(n)$ space. This computes for all $g\in V_{\sigma}$ their time interval $t_{1},t_{2}]$. Let $g\in V_{\sigma}$, $(g,[u_{\text{max}},v_{\text{max}}])\in\mathcal{B}_{\sigma}$ and $t_{1},t_{2}]$ be the corresponding interval. We create a weighted rectangle $R_{g}:=[t_{1},t_{2}]\times[u_{\text{max}},v_{\text{max}}]$ in $\mathbb{R}^{2}$ where the weight is $v_{\text{max}}$. This creates a set $R$ of $O(n)$ weighted rectangles in the plane, which we store in a stabbing-query data structure that for any query point $q$, returns the maximum-weighted rectangle in $R$ that intersects $q$ in logarithmic time. Such a stabbing query data structure of size $O(n)$ can be implemented in various ways through standard techniques (for details, we refer to the stabbing query implementations in [3]). Given a query point $x\in\partial P$ with $(i,j)\in\sigma$, we compute in $O(\log n)$ the corresponding time $t$ such that $(i,j)=(i_{t},j_{t})$. We then perform a stabbing query with the point $(t,x)$ and the maximum-weight rectangle corresponds to the desired query output. ∎
@@ 9  An $O(n\log n)$-time algorithm
We first check in $O(n\log n)$ time via Corollary˜24, if there is a solution of size at most $3$. Otherwise, via Theorem˜36, we compute the representation of next as a piecewise Möbius transform, consisting of $O(n)$ pieces. Given this representation, there is a $O(n\log k)$ algorithm [4], also discussed in [38, Remark 10], which can compute the minimum $k$, together with a point $x\in\partial P$, such that $\textnormal{{next}}^{k}(x)$ overtakes $x$, i.e., $x,\textnormal{{next}}^{k}(x)]=\partial P$, or conversely, $\textnormal{{next}}^{k}(x)\geq x+n$. From this $x$, via the representation of next we can compute a solution of size $k$ in time $O(k\log n)$, concluding the proof. ∎
@@ 10  Lower bound
Although this follows almost-immediately from the entropy argument in  [45], we present an argument for completeness. Consider the special case where for the input instance all elements in the list $A\cup B$ are distinct. Any deterministic correct algorithm must produce a certificate that verifies that for all index pairs $(i,j)$: $A[i]\neq B[j]$. Indeed, consider an algorithm $\mathcal{A}$ where for some input instance $I$, for some index pair $(i,j)$, the algorithm does not verify whether $A[i]=B[j]$. Then an adversary may simply consider two instances, $I$ and $I^{\prime}$ where $I^{\prime}$ replaces $B[j]$ by $A[i]$ – adding a duplicate. The algorithm $\mathcal{A}$ cannot distinguish between these two inputs, since any such distinction would produce a certificate that $A[i]\neq B[j]$ for the input $I$. However, this implies that the algorithm $\mathcal{A}$ must provide the wrong answer for one of these two inputs.
@@ 11  Conclusion
Another direction concerns the model of computation. Throughout this work we have ultilised the realRAM model, which is standard in computational geometry. One may ask how our results behave under different models of computation. For the classical Art Gallery problem, Abrahamsen, Adamaszek, and Miltzow [1] showed that even when polygon vertices are integers, an optimal solution may require guard positions with irrational coordinates and the realRAM is therefore arguably a necessary model of computation. The Contiguous Art Gallery problem does not exhibit this phenomenon. In fact, the existing solutions [9, 38] imply a mild upper bound on bit complexity: if $P$ is described using $\log n$ bits per coordinate, then any optimal solution requires at most $\tilde{O}(kn)$ bits. A formal analysis is given in [38].

### v2

@@ 3.2  Lower bound
In Section˜10, we reduce from Set Disjointness, where we are given two sets $A,B\subset[0,n^{3}]$, each of size $n$, and must decide whether $A\cap B=\emptyset$. While in the wordRAM-model there is a deterministic $O(n\log\log n)$-time algorithm for solving this problem (via sorting [Han2002Deterministic]), in comparison-based models of computation such as the realRAM-model, it has a worst-case lower bound of $\Omega(n\log n)$, even if $A$ is already sorted [YaoLowerBound]. Given $A,B\subset[1,n^{3}]$ with $A$ sorted, we can in linear time construct a simple polygon $P$ such that $A\cap B=\emptyset$ if and only if $P$ can be guarded with $2|B|+2$ guards (see Figure˜2).
@@ 4  A combinatorial classification of guards
We will define three collections of guards called dominators: the good, the bad, and the ugly. We prove that all good guards are dominated by a good dominator, which will induce a set $X$ of $O(n)$ points along $\partial P$ such that there exists an optimal solution that includes a guard $(g,[x,v])$ for $x\in X$. We prove that all bad guards are dominated by either a bad or an ugly dominator. The flowchart in Figure 3 illustrates our approach. To this end, we first make some observations:
@@ 4.3  The next function and our high-level algorithm
Definition 8 (The next function [theEnemy, RobsonSpaldingZheng2024_AnalyticArcCover]).
@@ 5  Intermezzo: data structures and sliding sequences
Data Structure 1 ([Guibas1987Optimal]).
@@
The data structure of [Guibas1987Optimal] stores a family of shortest paths in $P$ as balanced binary trees. For any query points $s,t\in P$, the shortest path $S(s,t)$ can be expressed as $O(\log n)$ subtrees, together with $O(\log n)$ newly computed edges called bridges.
The second property follows from a standard adaptation of this data structure (see also [Eades2020Visibility]): for each pair of consecutive edges on a precomputed shortest path, determine whether the turn is left or right. For each node in the tree, store a symbol indicating whether all of its descendant edges are left-turning, right-turning, or mixed. For a query, we can determine for each of the $O(\log n)$ newly computed bridges whether the turn at its endpoints is left or right. Combined with the pre-stored symbols, this information suffices to answer the second query. ∎
Data Structure 2 ([Hershberger1993Ray]).
@@
Throughout the paper, we assume that we have access to ˜1 and ˜2 for the given polygon. We repeatedly use the concept of a sliding sequence. A sliding sequence can be thought of as a pair of indices $(i,j)$ that can be updated by incrementing either $i$ or $j$. Formally we define in Definition˜1 a conforming sliding sequence used by almost all of our algorithms. For each $(i,j)$, we maintain the visibility core $\mathcal{E}[i-1,j+1]$ as a convex-hull data structure supporting emptiness, containment, extreme-point, and ray-shooting queries, and an additional angle query (defined later). For any linear-size sliding sequence $\{(i_{t},j_{t})\}$, Chan, Hershberger, and Pratt [Chan2019Time-Windowed] give a data structure supporting such updates and queries in $O(\log n)$ time:
Lemma 9 (Paraphrased Lemma 1 in [Chan2019Time-Windowed]).
@@
The above two lemmas imply the following, see also the remarks made in [Chan2019Time-Windowed].
@@
Let $(i,j)$ be the dynamic index pair from the Sliding-Window data structure (3). We maintain $\mathcal{E}[i-1,j+1]$ and $\mathcal{E}[i-1,j+2]$ separately, subject to increments of $i$ and $j$ via Corollary˜11. This maintains the vertices of these convex areas as balanced binary trees where the left-to-right traversal of the leaves corresponds to a clockwise traversal. For a a convex polygon with $n$ vertices, whose vertices are stored in a balanced tree according to their cyclic order, the containment, ray shooting, and extreme-point queries can be answered in $O(\log n)$ time [DOBKIN1983]. We can thus immediately support all queries except for angle queries.
@@ 6.3  A linear-size set $\mathcal{B}_{\sigma}$ that contains all reduced good dominators in $\mathcal{B}$
We store the active points in the data structure of Brodal and Jacob [brodal2002dynamic], which supports insertions and deletions in $O(\log n)$ time and and extreme-point queries in $O(\log n)$ time. Each time we add a vertex to the active set, we perform an insertion. When $j^{\prime}$ is incremented, we let $H$ be the half-plane left of $j^{\prime}+1,j^{\prime}+2]$ and iteratively perform extreme-point queries in the direction of the outward normal of $H$. If a query returns a vertex $x$ outside $H$, then we set $r_{x}=[j^{\prime}+1,j^{\prime}+2]$ and remove $x$ from the active set, charging $O(\log n)$ time for both the query and the update. The first time a query returns a vertex inside $H$, all remaining active points lie in $H$, and the process terminates.
@@
To construct the corresponding querying data structure, we loop over all $(i,j)\in\sigma$ in order, which discretizes time $t\in[0,|\sigma|]$ where at time $t$ our loop is at $(i_{t},j_{t})\in\sigma$. For each $g\in V_{\sigma}$ there is a unique time interval $t_{1},t_{2}]$ where for all $t\in[t_{1},t_{2}]$, $g$ is a vertex of $\mathcal{E}[i_{t}-1,j_{t}+1]$. We loop over all $t$ and use the first-in-first-out convex hull data structure from Corollary˜11 to maintain $\mathcal{E}[i_{t}-1,j_{t}+1]$ explicitly in $O(n\log n)$ time and $O(n)$ space. This computes for all $g\in V_{\sigma}$ their time interval $t_{1},t_{2}]$. Let $g\in V_{\sigma}$, $(g,[u_{\text{max}},v_{\text{max}}])\in\mathcal{B}_{\sigma}$ and $t_{1},t_{2}]$ be the corresponding interval. We create a weighted rectangle $R_{g}:=[t_{1},t_{2}]\times[u_{\text{max}},v_{\text{max}}]$ in $\mathbb{R}^{2}$ where the weight is $v_{\text{max}}$. This creates a set $R$ of $O(n)$ weighted rectangles in the plane, which we store in a stabbing-query data structure that for any query point $q$, returns the maximum-weighted rectangle in $R$ that intersects $q$ in logarithmic time. Such a stabbing query data structure of size $O(n)$ can be implemented in various ways through standard techniques (for details, we refer to the stabbing query implementations in [Agarwal2005Stabbing]). Given a query point $x\in\partial P$ with $(i,j)\in\sigma$, we compute in $O(\log n)$ the corresponding time $t$ such that $(i,j)=(i_{t},j_{t})$. We then perform a stabbing query with the point $(t,x)$ and the maximum-weight rectangle corresponds to the desired query output. ∎
@@ 9  An $O(n\log n)$-time algorithm
We first check in $O(n\log n)$ time via Corollary˜24, if there is a solution of size at most $3$. Otherwise, via Theorem˜36, we compute the representation of next as a piecewise Möbius transform, consisting of $O(n)$ pieces. Given this representation, there is a $O(n\log k)$ algorithm [aggarwal1989finding], also discussed in [RobsonSpaldingZheng2024_AnalyticArcCover, Remark 10], which can compute the minimum $k$, together with a point $x\in\partial P$, such that $\textnormal{{next}}^{k}(x)$ overtakes $x$, i.e., $x,\textnormal{{next}}^{k}(x)]=\partial P$, or conversely, $\textnormal{{next}}^{k}(x)\geq x+n$. From this $x$, via the representation of next we can compute a solution of size $k$ in time $O(k\log n)$, concluding the proof. ∎
@@ 10  Lower bound
Although this follows almost-immediately from the entropy argument in  [YaoLowerBound], we present an argument for completeness. Consider the special case where for the input instance all elements in the list $A\cup B$ are distinct. Any deterministic correct algorithm must produce a certificate that verifies that for all index pairs $(i,j)$: $A[i]\neq B[j]$. Indeed, consider an algorithm $\mathcal{A}$ where for some input instance $I$, for some index pair $(i,j)$, the algorithm does not verify whether $A[i]=B[j]$. Then an adversary may simply consider two instances, $I$ and $I^{\prime}$ where $I^{\prime}$ replaces $B[j]$ by $A[i]$ – adding a duplicate. The algorithm $\mathcal{A}$ cannot distinguish between these two inputs, since any such distinction would produce a certificate that $A[i]\neq B[j]$ for the input $I$. However, this implies that the algorithm $\mathcal{A}$ must provide the wrong answer for one of these two inputs.
@@ 11  Conclusion
Another direction concerns the model of computation. Throughout this work we have ultilised the realRAM model, which is standard in computational geometry. One may ask how our results behave under different models of computation. For the classical Art Gallery problem, Abrahamsen, Adamaszek, and Miltzow [abrahamsen_et_al:LIPIcs.SoCG.2017.3] showed that even when polygon vertices are integers, an optimal solution may require guard positions with irrational coordinates and the realRAM is therefore arguably a necessary model of computation. The Contiguous Art Gallery problem does not exhibit this phenomenon. In fact, the existing solutions [theEnemy, RobsonSpaldingZheng2024_AnalyticArcCover] imply a mild upper bound on bit complexity: if $P$ is described using $\log n$ bits per coordinate, then any optimal solution requires at most $\tilde{O}(kn)$ bits. A formal analysis is given in [RobsonSpaldingZheng2024_AnalyticArcCover].